A student has completed 72 credits at a college. 18 of those credits were A’s, 33 were B’s,
15 were C’s, and 6 were D’s. If an A is worth 4 points, a B 3 points, a C 2 points, and a D 1 point, what is the student’s GPA? Round answer to the hundredth place.

Answer by jorel1380(3719)

(18x4)+33(3)+15(2)+6(1) /72=207/72=2.875 weighted GPA
☺☺☺☺